Grade 11General Physics

find vector sums of forces A and B. Magnitude of A is 130N and is along OA where A is (3,-4,12) and magnitude of B is 200N and is along vector PQ where P is (-2,4,7) and Q is (-2,8,10)

Kshitij Ingle
10 Years agoGrade 11

3 Answers

thatikonda
10 Years ago
please”approve”............................................................................................................................................................................................................................................................................................
Kishorkumar
ApprovedApproved Tutor Answer10 Years ago
Given Vector magnitude of vector A=130 N
                      magnitude of vector B=200 N
     also stated that A is along vector OA (where O is orign & A(3,-4,12))
               & B is along the vector PQ where P(-2,4,7) Q(-2,8,10)
First we find the Unit vectors for both OA and PQ
Unit veector of OA=vector OA divided by Magnitude of OA
                               =3i-4j+12k /  Sq.Root of 32+(-4)2+122
                               =3/13 i – 4/13 j +12/13 k
 
Now Unit vector of PQ=vector PQ divided by Magnitude of PQ
                                     ={ -2-(-2)i+8-(4)j +10-(7)k} /  Sq.Root of 02+42+32
                                     =0i+4j+3k / 5
                                     =4/5 j +3/5 k
 
Therefore Vector A={3/13 i -4/13 j +12/13 k}×130
                              =30i-40j+120k
                 Vector B={4/5 i+ 3/5 k}×200
                               =160i +120k
Vector Sum of A and B={ (0-30)i+(160+40)j+(120-120)k}
                                    =-30i+200j
Hence Solved.
Koyal Chauhan
5 Years ago
Given Vector magnitude of vector A=130 N
                      magnitude of vector B=200 N
     also stated that A is along vector OA (where O is orign & A(3,-4,12))
               & B is along the vector PQ where P(-2,4,7) Q(-2,8,10)
First we find the Unit vectors for both OA and PQ
Unit veector of OA=vector OA divided by Magnitude of OA
                               =3i-4j+12k /  Sq.Root of 32+(-4)2+122
                               =3/13 i – 4/13 j +12/13 k
 
Now Unit vector of PQ=vector PQ divided by Magnitude of PQ
                                     ={ -2-(-2)i+8-(4)j +10-(7)k} / Sq.Root                                              of 02+42+32
                                     =0i+4j+3k / 5
                                     =4/5 j +3/5 k
 
Therefore Vector A={3/13 i -4/13 j +12/13 k}×130
                              =30i-40j+120k
                 Vector B={4/5 i+ 3/5 k}×200
                               =160i +120k
Vector Sum of A and B={ (0+30)i+(160+40)j+(120+120)k
                                        =30i+120j+240k